# Generalized Bayesian Multidimensional Scaling and Model Comparison

### v2

@@ 4 Simulation Studies
The primary objective of the simulation studies is to evaluate the performance of various models under diverse data structures. We compared several candidate models, denoted as $\mathcal{M}_{g}^{\mathcal{D}}$, where $g$ represents the model distribution for dissimilarities and $\mathcal{D}$ is the dissimilarity metric used during estimation. We examined three experimental settings: one with known underlying dimensions, one with skewed errors, and the other with outliers. We also conducted a runtime analysis to compare the computation costs of various candidate models using both the proposed annealed SMC algorithm and the MCMC algorithm introduced in Oh and Raftery (2001). To maintain comparability between the two Bayesian methods, we kept a consistent computational budget. Initially, the annealed SMC was executed with $K$ particles, and the number of iterations was recorded. Subsequently, we equitably allocated this computational budget to MCMC by setting the number of MCMC iterations equal to the product of the annealed SMC iterations and the number of particles. For the implementation of the MCMC algorithm in Oh and Raftery (2001), we utilized the available R package BayMDS by Oh and Lee (2022), which implements the truncated Gaussian model.
We established the values for the prior parameters by utilizing empirical Bayes methods, following the recommendations outlined in Oh and Raftery (2001). For the prior of $\sigma^{2}$, we chose $a=5$ and $b=SSR/m$ obtained from CMDS. For the prior of $\psi$, we chose $c=-2$ and $d=2$. For the hyperprior of $\lambda_{k}$, we set $\alpha=1/2$ and $\beta_{k}=\frac{1}{2}v_{k}^{(0)}/n$, where $v_{k}^{(0)}/n$ is the sample variance of the $k$th coordinate of X from CMDS. For the mixing distribution of $\zeta_{i,j}$, we used degrees of freedom $\nu=5$. These parameter values deliver satisfactory results in the simulation studies, and the same values of the prior parameters are used in all examples unless otherwise specified.
In the random walk Metropolis-Hastings step, the constant multiplier of the variance of the Gaussian proposal density for generating $\textbf{x}_{i}$ and $\sigma^{2}$ is chosen based on the characteristics of the data to ensure rapid mixing. Since the number of significant attributes $p$ is often unknown, most examples use $p=2$ for the purpose of visualization. In the annealed SMC algorithm, we set the number of particles to $K=200$, the rCESS threshold to $\phi=0.8$, and the resampling threshold to $\epsilon=0.5$. Unless otherwise specified in the example, all parameters are sampled during the particle propagation steps. Parallel computation is also implemented for the annealed SMC algorithm with 3 cores across all simulations and examples.
@@ 4.1 Experiment 1: Data with known true underlying dimension
where the dissimilarity metric $\mathcal{D}$ is chosen to be the Euclidean distance. Given $\delta_{i,j}$, the observed dissimilarities $d_{i,j}$ are generated from a normal distribution with mean $\delta_{i,j}$ and standard deviation of 1. These dissimilarities are truncated at 0. Thus, the simulated data consists of a 100 $\times$ 100 symmetric matrix of dissimilarities calculated from Euclidean distances with Gaussian errors. We included one frequentist method, CMDS, along with two Bayesian methods, MCMC and annealed SMC (ASMC), in the performance evaluation. In ASMC, all three candidate models were considered, and both STRESS values and log marginal likelihood estimates (logM) were reported.
@@ 4.2 Experiment 2: Data with skewed errors
In this example, we test how the proposed model performs when data skewness is present. We generated a dataset consisting of $n$ accurate/unobserved observations $\textbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$. Each of the observations has a dimension of 20, i.e., $\mathbf{x}_{i}=(x_{i,1},\ldots,x_{i,20})^{\top}$. The observations were simulated as a mixture of $50\%$ from $\mathcal{N}(0,1)$, $25\%$ from $\mathcal{N}(100,10)$ and $25\%$ from $\mathcal{N}(-10,1)$. Next, we generated $n$ noisy/observed observations Z of size 20 through a two-step process. First, we introduced minor errors into all observations X with the aim of simulating the systematic errors arising from data measurement. Second, varied percentages of the observations are subject to the contamination of moderate and significant errors, with the intention of replicating the scenario in which some observations are inaccurately recorded during data measurement. Specifically, moderate and significant errors were introduced into 20% and 2% of the observations, respectively. The Euclidean metric was then applied to obtain the dissimilarities $d_{i,j}$’s from the noisy observations Z and the dissimilarities $\tilde{d}_{i,j}$’s from accurate observations X. The errors $\epsilon_{i,j}$’s were computed by:
$\epsilon_{i,j}=d_{i,j}-\tilde{d}_{i,j},\qquad i\neq j,\;i,j=1,\ldots,n.$
A detailed description of the data generation process with skewed errors is given in Supplementary. We conducted experiments in several scenarios with sample sizes of $n=100,300,500,700$ to assess the scalability of the model. Each scenario is repeated for 20 runs. A histogram of the errors $\epsilon_{i,j}$’s from one run with $n=300$ is shown in Figure LABEL:sub@fig:sim2_me_euc. Our goal is to compare the performance of the proposed model $\mathcal{M}_{TSN}^{\text{Euclidean}}$ with the standard model with truncated Gaussian $\mathcal{M}_{TN}^{\text{Euclidean}}$ using the log marginal likelihoods when data are contaminated by some skewed errors.
@@
Figure LABEL:sub@fig:sim2_runtime depicts the comparison of the computation times, showing the advantages of the annealed SMC algorithm over the MCMC algorithm across all scenarios with varying sample sizes. We examined four scenarios to compare the performance of the annealed SMC algorithm and the MCMC algorithm under the identical truncated Gaussian model. This comparison helped illustrate how computation time varies with increasing sample size. The computation times typically show greater variability as the sample size increases, as observed for $n=100,300$, and $500$. In the case of $n=700$, the narrower boxes are attributed to some incomplete MCMC runs caused by out-of-memory issues. Therefore, we had to restrict the results to successful runs with fewer ASMC and MCMC iterations. These results illustrate that, even with more complicated models such as the truncated skewed Gaussian, our runtime remains comparable and even faster, demonstrating the scalability of the proposed framework to larger datasets. Figure LABEL:sub@fig:sim2_euclidean_logZ shows the performance comparison in terms of the log marginal likelihood as the skewed error presents. The findings of the study demonstrate that the model incorporating a truncated skewed Gaussian exhibits better performance, as the data under consideration is primarily with skewed errors. The results provide evidence that the skewed distributions are necessary for certain circumstances for modeling purposes.
@@ 5 Data Applications
5.1 NIPS text data with incremental dimensions
In the first example, we demonstrate the performance of the adaptive inference with the annealed SMC algorithm on text data with incremental dimensions. The text data is generated from articles from the Conference on Neural Information Processing Systems (NIPS). The NIPS dataset contains NIPS conference papers published between 1987 and 2015 (Dua and Graff, 2017; Perrone et al., 2017). In this study, our focus is directed toward a subset of the NIPS dataset, comprising a matrix of word counts extracted from 55 articles. This matrix is referred to as the document-term matrix, which is constructed after tokenization, stop-word removal, and truncation of the corpus by retaining only words that appear more than fifty times. The document-term matrix has counts for a list of 15005 words. Instead of Euclidean dissimilarity, we consider Cosine dissimilarity, which is suitable for discrete data such as word counts since it measures how dissimilar the documents are irrespective of their sizes.
We fitted the model $\mathcal{M}^{\text{Cosine}}_{TT}$ and compared the results in terms of STRESS values and computational times. The number of significant attributes $p$ is assumed to be 2. In this toy example, we considered the cases with $n_{0}=10,50$, and $n_{1}=1,5$. For each combination of $n_{0}$ and $n_{1}$, we considered two cases: one applies the annealed SMC algorithm with a fixed dimension to all $n=n_{0}+n_{1}$ observations, while the other employs the annealed SMC algorithm with an incremental dimension, leveraging the results from the first $n_{0}$ observations are already known. We set the number of particles to $K=100$ and the rCESS threshold to $\phi=0.9$ in this example. At each particle propagation step in the annealed SMC algorithm, a subset of two parameters is randomly selected for updating. If the selected parameter includes $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, a random subset containing half of its components is updated. This setting balances computational efficiency with effective exploration of the parameter space.
From Table 2, it can be seen that the STRESS values from both cases are close, and this validates the performance of the annealed SMC algorithm for incremental dimension. For the two cases with $n_{0}=10$, the computation times decrease by an average of 24% when applying the annealed SMC algorithm for incremental dimension. Similarly, for the two cases with $n_{0}=50$, the computation times fall by 31%. Moreover, the variability in computation times decreased in both cases with the implementation of adaptive inference.
Observations STRESS
Time
(in sec)
Observations STRESS
Time
(in sec)
n = 11 0.7093 19.0 (4.2) n = 51 0.7449 143.0 (36.3)
$n_{0}=10,n_{1}=1$ 0.7133 15.1 (3.1) $n_{0}=50,n_{1}=1$ 0.7419 100.9 (13.9)
n = 15 0.7194 29.4 (9.9) n = 55 0.7545 156.5 (37.0)
$n_{0}=10,n_{1}=5$ 0.6930 21.6 (4.0) $n_{0}=50,n_{1}=5$ 0.7536 104.4 (14.4)
5.2 Geographical data
The second example aims to study the performance of the proposed method and present visualizations of the estimations with uncertainty measures. As an illustrative example to study the performances of the CMDS and BMDS methods, we considered the US cities dataset from the US Census Bureau (Census Bureau, 2021), which contains Latitude and Longitude information from 15 large US cities. To evaluate the performances of the GBMDS, we appended 10 noise variables to add complexity. These noise variables are generated from a Gaussian distribution with a mean of 0 and variance comparable to the “Latitude” or “Longitude” variables.
We performed experiments across three scenarios, each of which had a distinct set of noises added to the data. The noises are incorporated into the data by assigning varying weights to the true and noisy variables. We represent the signal-to-noise ratio as $R_{\text{s:n}}$, which is defined as the ratio of the weight of the true signal to that of the noise. If $R_{\text{s:n}}>1$, it indicates that there is more signal than noise. In the first scenario, we assigned equal importance to all variables, resulting in $R_{\text{s:n}}=1$. In this case, the results depend equally on the signal and noisy variables. In the second scenario, we placed more emphasis on crucial variables such as “Latitude” and “Longitude” by decreasing the weights assigned to redundant noises, setting $R_{\text{s:n}}=4$. In the third scenario, we tested an extreme condition where the majority of the weights were allocated to the “Latitude” and “Longitude” variables, setting $R_{\text{s:n}}=10$. In all experiments, we normalized the weights to ensure they sum up to 1.
We employed two Bayesian methods, MCMC and annealed SMC (ASMC), to implement our proposed GBMDS. To initialize the GBMDS, we utilized the results from CMDS. For simplicity, we assumed $\psi=0$ to reduce the model to $\mathcal{M}_{TN}^{\text{Euclidean}}$. To ensure a fair comparison between the two Bayesian methods, we kept the computational budget constant. Specifically, we first ran annealed SMC with 300 particles and recorded the number of iterations. We then allocated the same budget to MCMC by setting the number of MCMC iterations equal to the product of the annealed SMC iterations and the number of particles. The purpose of this example is to highlight the advantages of the Bayesian method over the frequentist method. For small datasets under the same model $\mathcal{M}_{TN}^{\text{Euclidean}}$, it is expected that the two Bayesian methods would provide similar results and MDS reconstructions of city locations, provided we run sufficiently long MCMC iterations in MCMC and use a large number of particles in ASMC. Therefore, we excluded the results from MCMC as both Bayesian methods yield similar results on this small dataset. This exclusion does not affect the overall conclusions, but rather ensures a clearer presentation of the key findings.
Table 3 presents the STRESS values obtained from CMDS and GBMDS with ASMC (GBMDS-ASMC) for the three scenarios. Our analysis indicates that the Bayesian approach yields lower STRESS values across all three scenarios. Figure 4 displays the estimated locations of the 15 US cities obtained by GBMDS-ASMC. Transformations, such as rotation and reflection, were applied to the estimated locations from GBMDS-ASMC to fit the cities’ actual geographical locations. One can observe from Figures LABEL:sub@fig:ex2_cmds_R_1 and LABEL:sub@fig:ex2_bmds_R_1 that under the equal weight scenario, several cities are geographically misplaced no matter what transformations are applied. The reason behind this mismatch is that the information in the “Latitude” and “Longitude” variables are masked by the remaining variables. The estimated locations shown in LABEL:sub@fig:ex2_cmds_R_10 and LABEL:sub@fig:ex2_bmds_R_10 lead to a closer match when higher weights are assigned to the “Latitude” and “Longitude” variables.
CMDS GBMDS-ASMC
Scenario 1: $R_{\text{s:n}}=1$ 0.4557 0.3521
Scenario 2: $R_{\text{s:n}}=4$ 0.4680 0.3910
Scenario 3: $R_{\text{s:n}}=10$ 0.4726 0.4103
The Bayesian approach offers several advantages over the classical approach. In addition to producing smaller STRESS values, the Bayesian approach enables the estimation of uncertainty by leveraging samples from the posterior distribution. To this end, we performed Procrustes transformations on the posterior samples of $\textbf{x}_{i}$ to align each sample as closely as possible to the estimated coordinates, effectively standardizing all the posterior samples of $\textbf{x}_{i}$. Using the transformed posterior samples, we constructed credible regions, represented as ellipses in Figures LABEL:sub@fig:ex2_bmds_R_1 to LABEL:sub@fig:ex2_bmds_R_10. In contrast to CMDS, our GBMDS-ASMC method offers uncertainty measures, with tight credible regions in scenarios where the signal-to-noise ratio is high and wider credible regions when more noises exist in the data.
@@ 5.3 NIH text data
In this example, we apply the MDS techniques to text data consisting of documents and words. We aim to showcase the application of our proposed method and investigate the effect of dimension $p$. The dataset holds information on research grants awarded by the National Institutes of Health (NIH) in 2022 (NIH, 2022). The raw data contain 1471 grant abstracts from Saint Louis in the United States. To preprocess the data, we performed tokenization and removed stop words. This results in a document-term matrix with a dimension of 1471 by 238919, where each row represents an abstract and contains the word counts for all words in that abstract. We then re-weighted the word counts by multiplying an inverse document frequency vector to adjust for the relative importance of words in the entire collection of documents. The purpose of this reweighing step was to account for the varying frequencies of words across documents. We calculated the Cosine dissimilarities of the documents and used them as input to the MDS methods. Figure 5a displays the histogram of Cosine dissimilarity values, with $Y$-axis on a log scale indicating the log count of occurrences. The histogram shows a spike at the lower end, around 0, and then a gradually increasing trend in counts as the Cosine dissimilarities approach 1. This distribution suggests that the dataset contains pairs with high similarity (values near 0) and high dissimilarity (values near 1), with a notable concentration at higher dissimilarity levels.
@@ 5.4 Clustering abstract text data
The last example explores an interesting application of MDS, clustering objects by grouping similar observations based on their dissimilarities. With MDS techniques, the coordinates of objects can be represented in a low-dimensional space. Visualizing these clusters in the reduced-dimensional space is of interest, as it can provide meaningful insights into the relationships between the groups. The dataset comprises 139 English abstracts collected from three academic journals during 2022 and 2023: 44 from Annals of Mathematics, 50 from Bayesian Analysis, and 45 from Biometrics. While all three journals focus on the development of statistical and mathematical methods, each emphasizes distinct application areas. This study aims to assess the performance of the proposed model in clustering these abstract texts. We employed the Jaccard metric with a 2-gram model to construct the dissimilarity matrix. A 2-gram model represents a subsequence of two successive elements within an abstract, including words, numbers, or symbols. Each abstract was first treated as a set of unique 2-grams, and the Jaccard dissimilarity was computed based on the intersection and union of these sets. This metric is effective for identifying abstracts with overlapping vocabularies or patterns, mainly when abstracts are dense with technical terms. Additionally, the use of a 2-gram model is beneficial in applications where sequences of words are relevant.
We applied GBMDS-ASMC for reducing dimension, followed by the partition around medoids algorithm (PAM) (Kaufman, 1990) to group the results into 3 clusters. In this study, we predetermined the number of clusters as 3, corresponding to the three journals included in the dataset. All candidate models were evaluated, and the model $\mathcal{M}^{\text{Euclidean}}_{TSN}$ at dimension $p=7$ achieved the highest log marginal likelihood estimates. Therefore, it was selected as the optimal model for this example. Figure 6 presents the pairwise scatterplots and density plots of the clustering results, highlighting that the first three dimensions contribute significantly to the clustering, with clear separations between clusters. This result provides a meaningful representation of the data, yielding well-defined clusters with visual separation. We also evaluated the classification accuracy as another performance metric using the known journal labels of each abstract. The GBMDS-ASMC method correctly classified all abstracts into their respective journals. Table 4 summarizes the top 20 keywords from each cluster along with their corresponding journals. The results reveal the presence of similar keywords related to the development of statistical and mathematical methods. Additionally, distinct keywords are observed, reflecting the specific application and thematic focus unique to each journal.

### v3

@@ 4 Simulation Studies
The primary objective of the simulation studies is to evaluate the performance of various models under diverse data structures. We compared several candidate models, denoted as $\mathcal{M}_{g}^{\mathcal{D}}$, where $g$ represents the model distribution for dissimilarities and $\mathcal{D}$ is the dissimilarity metric used during estimation. We examined three experimental settings: one with known underlying dimensions, one with skewed errors, and the other with outliers. We also conducted a runtime analysis to compare the computation costs of various candidate models using both the proposed annealed SMC algorithm and the MCMC algorithm introduced in oh2001bayesian. To maintain comparability between the two Bayesian methods, we kept a consistent computational budget. Initially, the annealed SMC was executed with $K$ particles, and the number of iterations was recorded. Subsequently, we equitably allocated this computational budget to MCMC by setting the number of MCMC iterations equal to the product of the annealed SMC iterations and the number of particles. For the implementation of the MCMC algorithm in oh2001bayesian, we utilized the available R package BayMDS by oh2022baymds, which implements the truncated Gaussian model.
We established the values for the prior parameters by utilizing empirical Bayes methods, following the recommendations outlined in oh2001bayesian. For the prior of $\sigma^{2}$, we chose $a=5$ and $b=SSR/m$ obtained from CMDS. For the prior of $\psi$, we chose c=2c=-2 and $d=2$. For the hyperprior of $\lambda_{k}$, we set $\alpha=1/2$ and $\beta_{k}=\frac{1}{2}v_{k}^{(0)}/n$, where $v_{k}^{(0)}/n$ is the sample variance of the $k$th coordinate of X from CMDS. For the mixing distribution of $\zeta_{i,j}$, we used degrees of freedom $\nu=5$. The variance parameter for the reference distribution of the initial latent variables $\textbf{x}_{i}$ is set to $\rho=0.01$. These parameter values deliver satisfactory results in the simulation studies, and the same values of the prior parameters are used in all examples unless otherwise specified.
In the random walk Metropolis-Hastings step, the constant multiplier of the variance of the Gaussian proposal density for generating $\textbf{x}_{i}$ and $\sigma^{2}$ is chosen based on the characteristics of the data to ensure rapid mixing. In the annealed SMC algorithm, we set the number of particles to $K=200$, the rCESS threshold to $\phi=0.8$, and the resampling threshold to $\epsilon=0.5$. Unless otherwise specified in the example, all parameters are sampled during the particle propagation steps. Parallel computation is also implemented for the annealed SMC algorithm with 3 cores across all simulations and examples.
@@ 4.1 Experiment 1: Data with known true underlying dimension
where the dissimilarity metric $\mathcal{D}$ is chosen to be the Euclidean distance. Given $\delta_{i,j}$, the observed dissimilarities $d_{i,j}$ are generated from a normal distribution with mean $\delta_{i,j}$ and standard deviation of 1. These dissimilarities are truncated at 0. Thus, the simulated data consists of a 100 $\times$ 100 symmetric matrix of dissimilarities calculated from Euclidean distances with Gaussian errors. We included one frequentist method, CMDS, along with two Bayesian methods, MCMC and annealed SMC (ASMC), in the performance evaluation. In ASMC, all three candidate models ($\mathcal{M}_{TN},\mathcal{M}_{TSN}$, and $\mathcal{M}_{TT}$) were considered, and both STRESS values and log marginal likelihood estimates (log $Z$) were reported.
@@ 4.2 Experiment 2: Data with skewed errors
In this example, our goal is to compare the performance of the proposed model $\mathcal{M}_{TSN}^{\text{Euclidean}}$ against the standard truncated Gaussian model $\mathcal{M}_{TN}^{\text{Euclidean}}$ in terms of log marginal likelihoods, particularly when the data are contaminated by skewed errors. We test how the proposed model performs when data skewness is present. We generated a dataset consisting of $n$ data points $\check{\textbf{Z}}=\left\{\check{\mathbf{z}}_{1},\ldots,\check{\mathbf{z}}_{n}\right\}$. Each of them has a dimension of 20, i.e., $\check{\mathbf{z}}_{i}=(\check{z}_{i,1},\ldots,\check{z}_{i,20})^{\top}$. The data were simulated as a mixture of $50\%$ from $\mathcal{N}(0,1)$, $25\%$ from $\mathcal{N}(100,10)$ and $25\%$ from $\mathcal{N}(-10,1)$. Next, we generated $n$ noisy observations Z through a two-step process. First, we introduced minor errors into all data $\check{\textbf{Z}}$ to simulate systematic errors arising from data measurement. Second, varied percentages of the data points are subject to the contamination of moderate and significant errors, with the intention of replicating the scenario in which some observations are inaccurately recorded during data measurement. Specifically, moderate and significant errors were introduced into 20% and 2% of the data points, respectively. The Euclidean metric was then applied to obtain the dissimilarities $d_{i,j}$’s from the noisy data Z and the dissimilarities $\check{d}_{i,j}$’s from clean data $\check{\textbf{Z}}$. The errors $\epsilon_{i,j}$’s were computed by:
$\epsilon_{i,j}=d_{i,j}-\check{d}_{i,j},\qquad i\neq j,\;i,j=1,\ldots,n.$
A detailed description of the data generation process with skewed errors is given in Supplementary. We conducted experiments in several scenarios with sample sizes of $n=100,300,500,700$ to assess the scalability of the model. Each scenario is repeated for 20 runs. A histogram of the errors $\epsilon_{i,j}$’s from one run with $n=300$ is shown in Figure LABEL:sub@fig:sim2_me_euc.
@@
Figure LABEL:sub@fig:sim2_runtime depicts the comparison of the computation times, showing the advantages of the annealed SMC algorithm over the MCMC algorithm across all scenarios with varying sample sizes. We examined four scenarios to compare the performance of the annealed SMC algorithm and the MCMC algorithm under the identical truncated Gaussian model. This comparison helped illustrate how computation time varies with increasing sample size. The computation times typically show greater variability as the sample size increases, as observed for $n=100,300$, and $500$. These results illustrate that, even with more complicated models such as the truncated skewed Gaussian, our runtime remains comparable and even faster, demonstrating the scalability of the proposed framework to larger datasets. Figure LABEL:sub@fig:sim2_euclidean_logZ shows the performance comparison in terms of the log marginal likelihood as the skewed error presents. The findings of the study demonstrate that the model incorporating a truncated skewed Gaussian exhibits better performance, as the data under consideration is primarily with skewed errors. The results provide evidence that the skewed distributions are necessary for certain circumstances for modeling purposes.
@@ 5.1 NIH text data
In this example, we apply the MDS techniques to text data consisting of documents and words. We aim to showcase the application of our proposed method and investigate the effect of dimension $p$. The dataset holds information on research grants awarded by the National Institutes of Health (NIH) in 2022 [NIHRePORTER]. The raw data contain 1471 grant abstracts from Saint Louis in the United States. To preprocess the data, we performed tokenization and removed stop words. This results in a document-term matrix with a dimension of 1471 by 238919, where each row represents an abstract and contains the word counts for all words in that abstract. We then re-weighted the word counts by multiplying an inverse document frequency vector to adjust for the relative importance of words in the entire collection of documents. The purpose of this reweighing step was to account for the varying frequencies of words across documents. We calculated the Cosine dissimilarities of the documents and used them as input to the MDS methods. Figure 4a displays the histogram of Cosine dissimilarity values, with $Y$-axis on a log scale indicating the log count of occurrences. The histogram shows a spike at the lower end, around 0, and then a gradually increasing trend in counts as the Cosine dissimilarities approach 1. This distribution suggests that the dataset contains pairs with high similarity (values near 0) and high dissimilarity (values near 1), with a notable concentration at higher dissimilarity levels.
@@ 5.2 Clustering abstract text data
This example explores an interesting application of MDS, clustering objects by grouping similar observations based on their dissimilarities. With MDS techniques, the coordinates of objects can be represented in a low-dimensional space. Visualizing these clusters in the reduced-dimensional space is of interest, as it can provide meaningful insights into the relationships between the groups. The dataset comprises 139 English abstracts collected from three academic journals during 2022 and 2023: 44 from Annals of Mathematics, 50 from Bayesian Analysis, and 45 from Biometrics. While all three journals focus on the development of statistical and mathematical methods, each emphasizes distinct application areas. This study aims to assess the performance of the proposed model in clustering these abstract texts. We employed the Jaccard metric with a 2-gram model to construct the dissimilarity matrix. A 2-gram model represents a subsequence of two successive elements within an abstract, including words, numbers, or symbols. Each abstract was first treated as a set of unique 2-grams, and the Jaccard dissimilarity was computed based on the intersection and union of these sets. This metric is effective for identifying abstracts with overlapping vocabularies or patterns, mainly when abstracts are dense with technical terms. Additionally, the use of a 2-gram model is beneficial in applications where sequences of words are relevant.
We applied GBMDS-ASMC for reducing dimension, followed by the partition around medoids algorithm (PAM) [kaufman1990partitioning] to group the results into 3 clusters. In this study, we predetermined the number of clusters as 3, corresponding to the three journals included in the dataset. All candidate models were evaluated, and the model $\mathcal{M}^{\text{Euclidean}}_{TSN}$ at dimension $p=7$ achieved the highest log marginal likelihood estimates. Therefore, it was selected as the optimal model for this example. Figure 5 presents the pairwise scatterplots and density plots of the clustering results, highlighting that the first three dimensions contribute significantly to the clustering, with clear separations between clusters. This result provides a meaningful representation of the data, yielding well-defined clusters with visual separation. We also evaluated the classification accuracy as another performance metric using the known journal labels of each abstract. The GBMDS-ASMC method correctly classified all abstracts into their respective journals. Table 2 summarizes the top 20 keywords from each cluster along with their corresponding journals. The results reveal the presence of similar keywords related to the development of statistical and mathematical methods. Additionally, distinct keywords are observed, reflecting the specific application and thematic focus unique to each journal.
@@
5.3 NeurIPS text data with incremental data updates
In the first example, we demonstrate the performance of the adaptive inference with the annealed SMC algorithm on text data with incremental size. The text data is generated from articles from the Conference on Neural Information Processing Systems (NeurIPS). The NeurIPS dataset contains NeurIPS conference papers published between 1987 and 2015 [Dua:2019, perrone2017poisson]. In this study, our focus is directed toward a subset of the NeurIPS dataset, comprising a matrix of word counts extracted from 55 articles. This matrix is referred to as the document-term matrix, which is constructed after tokenization, stop-word removal, and truncation of the corpus by retaining only words that appear more than fifty times. The document-term matrix has counts for a list of 15005 words. Instead of Euclidean dissimilarity, we consider Cosine dissimilarity, which is suitable for discrete data such as word counts since it measures how dissimilar the documents are irrespective of their sizes.
We fitted the model $\mathcal{M}^{\text{Cosine}}_{TT}$ and compared the results by STRESS values and computational times. The number of significant attributes $p$ is assumed to be 2. In this toy example, we considered the cases with $m_{0}=10,50$, and $m_{1}=1,5$. For each combination of $m_{0}$ and $m_{1}$, we considered two cases: one applies the annealed SMC algorithm to the combined $n=m_{0}+m_{1}$ observations from scratch, while the other employs the annealed SMC algorithm with sequential data updates, leveraging the posterior samples obtained from the first $m_{0}$ observations. We set the number of particles to $K=100$ and the rCESS threshold to $\phi=0.9$. At each particle propagation step, a subset of two parameters is randomly selected for updating. If the latent variables $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ are selected, a random subset containing half of the components is updated. This setting balances computational efficiency with effective exploration of the parameter space.
From Table 3, it can be seen that the STRESS values from both cases are close, and this validates the performance of the annealed SMC algorithm for incremental dimension. For the two cases with $m_{0}=10$, the computation times decrease by an average of 24% when applying the annealed SMC algorithm for incremental data. Similarly, for the two cases with $m_{0}=50$, the computation times fall by 31%. Moreover, the variability in computation times decreased in both cases with the implementation of adaptive inference.
Observations STRESS
Time
(in sec)
Observations STRESS
Time
(in sec)
n = 11 0.7093 19.0 (4.2) n = 51 0.7449 143.0 (36.3)
$m_{0}=10,m_{1}=1$ 0.7133 15.1 (3.1) $m_{0}=50,m_{1}=1$ 0.7419 100.9 (13.9)
n = 15 0.7194 29.4 (9.9) n = 55 0.7545 156.5 (37.0)
$m_{0}=10,m_{1}=5$ 0.6930 21.6 (4.0) $m_{0}=50,m_{1}=5$ 0.7536 104.4 (14.4)
5.4 Geographical data
This example aims to study the performance of the proposed method and visualize the estimations with uncertainty measures. As an illustrative example to compare CMDS and BMDS, we considered the US cities dataset from the US Census Bureau [UScensus], which contains Latitude and Longitude information for 15 large US cities. To evaluate the robustness of GBMDS, we appended 10 noise variables generated from a Gaussian distribution with a mean of 0 and variance comparable to the “Latitude” or “Longitude” variables.
We performed experiments across three scenarios, differing in the weight assigned to true signal versus noise. We define the signal-to-noise ratio as $R_{\text{s:n}}$ as the ratio of the total weight of the true signal variables to that of the noise. In Scenario 1 ($R_{\text{s:n}}=1$), the total weight of the two signal variables equals the total weight of the 10 noise variables. In Scenario 2 ($R_{\text{s:n}}=4$), emphasis is placed on the geographic variables. Scenario 3 ($R_{\text{s:n}}=10$) represents an extreme condition where the majority of the weight is allocated to the signal. In all experiments, weights were normalized to sum to 1.
CMDS GBMDS-ASMC
Scenario 1: $R_{\text{s:n}}=1$ 0.4557 0.3521
Scenario 2: $R_{\text{s:n}}=4$ 0.4680 0.3910
Scenario 3: $R_{\text{s:n}}=10$ 0.4726 0.4103
To implement the proposed GBMDS, we utilized the annealed SMC (ASMC) algorithm, initialized using results from CMDS. For simplicity, we assumed $\psi=0$, reducing the model to $\mathcal{M}_{TN}^{\text{Euclidean}}$. While standard MCMC methods yield similar results for this small dataset1, we report only the ASMC results to maintain focus on the comparison between Bayesian and frequentist approaches.
Table 4 presents the STRESS values obtained from CMDS and GBMDS (via ASMC) for the three scenarios. The Bayesian approach yields lower STRESS values across all cases. Figure 6 displays the estimated locations of the 15 US cities. Procrustes transformations (rotation and reflection) were applied to align the estimates with the actual map. As observed in Figures LABEL:sub@fig:ex2_cmds_R_1 and LABEL:sub@fig:ex2_bmds_R_1, under the equal weight scenario ($R_{\text{s:n}}=1$), several cities are geographically misplaced regardless of transformation, as the geographic signal is masked by the noise variables. However, the estimated locations shown in LABEL:sub@fig:ex2_cmds_R_10 and LABEL:sub@fig:ex2_bmds_R_10 show a much closer match when higher weights are assigned to the “Latitude” and “Longitude” variables.
The Bayesian approach offers several advantages over the classical approach. Beyond producing lower STRESS values, the Bayesian approach enables the estimation of uncertainty by leveraging samples from the posterior distribution. Since MDS solutions are invariant to rotation and translation, we performed Procrustes transformations on the posterior samples of $\textbf{x}_{i}$ to align them to a reference configuration (the posterior mean estimate), effectively standardizing the orientation of all samples of $\textbf{x}_{i}$. Using these aligned samples, we constructed credible regions, represented as ellipses in Figures LABEL:sub@fig:ex2_bmds_R_1 to LABEL:sub@fig:ex2_bmds_R_10. In contrast to CMDS, our GBMDS-ASMC method provides quantifiable uncertainty measures, exhibiting tight credible regions in scenarios with high signal-to-noise ratios and wider regions when the data contains significant noise.